OpenStudy (anonymous):

Hello everyone My cousin rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the STANDARD FORM to represent the total rent (y) that my cousin has to pay for renting the coat for x days.

OpenStudy (anonymous):

@angelrush101

OpenStudy (anonymous):

y = a +bx a is fixed cost, while b is variable cost so equation 1 & 2 are as follows, 285 = a +3b 510 = a + 6b can u solve for a & b ??

OpenStudy (anonymous):

Its a wonderful place called google XD

OpenStudy (anonymous):

Thank you @angelrush101 I got it from here.

OpenStudy (anonymous):

Actually @angelrush101 I need more help, please. I can't solve for a & b